# Improving Object Detection by Modifying Synthetic Data with Explainable AI

### v1

@@ 11 Angular breakdown of target misclassification and SHAP plots
At the end of Section 2.1 in the main paper, we note that the comparison of saliency maps to identify unique and common features between two classes must also account for variation across each class. In this dataset we have a natural dimension of variation (orientation of the vehicle), and so we do not need to cluster the samples to explore the dimension of variation. To identify which orientations of the target vehicle are being confused with the misclassified vehicle, we first provide a breakdown in segments of $5^{\circ}$ for the target misclassification between the SUV and BTR70 (e.g. the example shown in Fig. 5a-c in the main paper) and the misclassification from the ZSU23 to the BTR70 (e.g. see Fig. 5d-f in the main paper). By providing this angular breakdown we motivate the identification of how confusion between the two different vehicles was identified to be aligned with opposite orientations for the SUV at $\theta=70^{\circ}-75^{\circ}$ to the BTR70 at $\theta=285^{\circ}-290^{\circ}$ and how similar orientations for the ZSU23 at $\theta=105^{\circ}-110^{\circ}$ to the BTR70 at $\theta=125^{\circ}-130^{\circ}$, as described in the main paper.
@@
Figure 14: (a) Fraction of the misclassifications of the SUV as BTR70 for each $5^{\circ}$ over the test dataset. The blue line with squares corresponds to synthetic model v0 before the modification and the green line with circles corresponds to synthetic model vR, after the reinforcing modification. The pink dashed rectangle indicates the orientation range considered for the misclassifications. (b) The average SHAP plot for the SUV misclassifications to BTR70 at $70^{\circ}-75^{\circ}$. (c) Average SHAP plots for the correct classifications of BTR70 at all orientations. The regions of common features associated with misclassification are highlighted by pink rectangles in panels (b) and (c)
@@
Figure 15: (a) Fraction of the misclassifications of the ZSU23 as the BTR70 for each $5^{\circ}$ over the test dataset. The blue line with squares corresponds to synthetic model v0 before the modification and the green line with circles corresponds to synthetic model vD, after the disruptive modification. The orange dashed rectangle indicates the orientation range considered for the misclassifications. (b) Average SHAP plot for the ZSU23 at $105^{\circ}-110^{\circ}$. (c) Average SHAP plots for the BTR70 at all orientations. The regions of common features associated with misclassification are highlighted by orange rectangles in panels (b) and (c)

### v2

@@ 11 Angular breakdown of target misclassification and SHAP plots
At the end of Section 2 in the main paper, we note that the comparison of saliency maps to identify unique and common features between two classes must also account for variation across each class. In this dataset we have a natural dimension of variation (orientation of the vehicle), and so we do not need to cluster the samples to explore the dimension of variation. To identify which orientations of the target vehicle are being confused with the misclassified vehicle, we first provide a breakdown in segments of $5^{\circ}$ for the target misclassification between the SUV and BTR70 (e.g. the example shown in Fig. 5a-c in the main paper) and the misclassification from the ZSU23 to the BTR70 (e.g. see Fig. 5d-f in the main paper). By providing this angular breakdown we motivate the identification of how confusion between the two different vehicles was identified to be aligned with opposite orientations for the SUV at $\theta\in[70^{\circ},75^{\circ}]$ to the BTR70 at $\theta\in[285^{\circ},290^{\circ}]$ and how similar orientations for the ZSU23 at $\theta\in[105^{\circ},110^{\circ}]$ to the BTR70 at $\theta\in[125^{\circ},130^{\circ}]$, as described in the main paper.
@@
Figure 23: (a) Fraction of the misclassifications of the SUV as BTR70 for each $5^{\circ}$ over the test dataset. The blue line with squares corresponds to synthetic model v0 before the modification and the green line with circles corresponds to synthetic model vR, after the reinforcing modification. The pink dashed rectangle indicates the orientation range considered for the misclassifications. (b) The average SHAP plot for the SUV misclassifications to BTR70 at $70^{\circ}-75^{\circ}$. (c) Maximum correlations of SUV misclassification SHAP plot in panel (b) with BTR70 correct classifications at different orientations. This helps to determine which subset (of orientations) of BTR70 the SUV is being misclassified as (highlighted by the pink dashed rectangle).(d) Average SHAP plots for the correct classifications of BTR70 at all orientations. The regions of common features associated with misclassification are highlighted by pink rectangles in panels (b) and (c)
@@
Figure 24: (a) Fraction of the misclassifications of the ZSU23 as the BTR70 for each $5^{\circ}$ over the test dataset. The blue line with squares corresponds to synthetic model v0 before the modification and the green line with circles corresponds to synthetic model vD, after the disruptive modification. The orange dashed rectangle indicates the orientation range considered for the misclassifications. (b) Average SHAP plot for the ZSU23 at $105^{\circ}-110^{\circ}$. (c) Maximum correlations of ZSU23 misclassification SHAP plot in panel (b) with BTR70 correct classifications at different orientations. This helps to determine which subset (of orientations) of BTR70 the ZSU23 is being misclassified as (highlighted by the orange dashed rectangle). (d) Average SHAP plots for the BTR70 at all orientations. The regions of common features associated with misclassification are highlighted by orange rectangles in panels (b) and (c)